Question 55: Please refer the scenario below for questions 55-56: All workers at one of Cannondale's factories completed a survey that asked if they were willing to ride a bicycle to work, should Cannondale provide them one for free. Of the 96 workers that lived less than five miles away, one third said they would. Of the 48 that lived between 5 and 10 miles away, 25% said they would. Of the 25 that lived between 10 and 15 miles away, 20% said that they would. Of the 20 between 15 to 20 miles away, only one workers said he would ride a bicycle to work if given one. There were no workers living more than 20 miles away. What was the total number of workers living within 15 miles that were NOT willing to ride a bicycle to work?

  1. 79
  2. 189
  3. 50
  4. 120

Question 56: What is the estimated average distance that workers live from the factory?

  1. 139 miles
  2. 6.68 miles

      C.    9.18 miles

      D.   18.9 miles
Answer #1

55) Total workers living within 15 miles = 96+48+25 = 169

Total workers willing to ride bicycle within 15 miles = (1/3)*96 + 0.25*48 + 0.2*25 = 49

Total workers not willing to ride bicycle within 15 miles = 169-49 = 120

So D is the correct answer

56) We will use the midpoint of the ranges given for the calculation of the mean distance at which the workers live.

So the average distance = (96*2.5 + 48*7.5 + 25*12.5 + 17.5*20) / (96+48+25+20)

= 1262.5 / 189

= 6.68 miles

So option B is the correct answer.